# MD-PK: Metaphor Detection via Prompt Learning and Knowledge Distillation

### v1

@@ Abstract
Metaphors are ubiquitous in daily life, yet detecting them poses a significant challenge. Previous approaches often struggled with improper application of language rules and overlooked the issue of data sparsity. To address these challenges, we introduce knowledge distillation and prompt learning into metaphor detection. Specifically, we devise a prompt learning template tailored for the metaphor detection task. By masking target words and providing relevant prompt information, we guide the model to accurately infer the contextual meaning of these words. This approach not only mitigates the interference from the literal meaning of target words but also ensures the proper utilization of MIP language rules for metaphor detection. Moreover, we employ a teacher model equipped with prior knowledge to generate meaningful soft labels, guiding the optimization process of the student model. The inclusion of soft labels, akin to label smoothing, helps alleviate the model’s tendency towards over-confidence and effectively addresses the challenge of data sparsity. Experimental results demonstrate that our proposed model achieves state-of-the-art performance across multiple datasets.
MD-PK: Metaphor Detection via Prompt Learning and Knowledge Distillation
@@ 1 Introduction
The task of metaphor detection poses significant challenges. Early methods Turney et al. (2011); Broadwell et al. (2013); Tsvetkov et al. (2014); Bulat et al. (2017) relied on hand-designed linguistic features to identify metaphors, while later approaches Wu et al. (2018); Gao et al. (2018); Mao et al. (2019) utilized recurrent neural networks (RNNs) to analyze contextual information. However, both categories heavily depended on manually crafted structures, and the encoders employed struggled to handle complex contextual information, leading to limited effectiveness. In contrast, more recent methods Gong et al. (2020); Su et al. (2020); Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023) leverage pre-trained language models such as BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019) to process contextual information, resulting in significantly improved performance.
While utilizing pre-trained language models has yielded the best results in metaphor detection, these approaches often fail to fully leverage linguistic rules. Among the most advanced methods Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023), the incorporation of linguistic rules, such as the Metaphor Identification Procedure (MIP) Group (2007), stands out. MIP operates on the principle of extracting the contextual meaning of a target word in the original sentence. If this contextual meaning diverges from the literal meaning, a metaphorical use is identified. For instance, in the sentence "We must bridge the gap between employees and management," the contextual meaning of the target word ’bridge’ pertains to reducing differences or divisions, while its literal meaning refers to constructing a physical bridge. The inconsistency between these meanings indicates a metaphorical use. However, despite the logical validity of MIP, current methods fail to effectively utilize it. Typically, these methods Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023) encode the sentence directly using an encoder and extract the word vector from the position of the target word to represent its contextual meaning. Yet, this approach is flawed as the word vector may be influenced by the literal meaning of the target word, leading to an incomplete representation of its contextual meaning.
@@ 3.2 Metaphor Detection
In this paper, we leverage two linguistic rules to aid in metaphor detection: Selectional Preference Violation (SPV) Wilks (1978) and Metaphor Identification Procedure (MIP) Group (2007). SPV operates on the principle of identifying metaphors by discerning inconsistencies between the target word and its context. Conversely, MIP detects metaphors by comparing the contextual meaning of the target word with its literal meaning.
When employing SPV linguistic rules for metaphor detection, we adopt a methodology similar to previous works Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023). Specifically, we encode the sentence to extract the target word vector and the sentence vector. Subsequently, we analyze the relationship between these vectors using a fully connected matrix to detect inconsistencies between the target word and its context.
@@ 3.3 Knowledge Distillation
Our proposed knowledge distillation method comprises a teacher model T and a student model S. The teacher model T is pre-trained with extensive prior knowledge, enabling it to generate meaningful soft labels to facilitate the optimization process of the student model. By learning from these soft labels generated by the teacher model, the student model S rapidly enhances its capabilities. Furthermore, given the prevalent issue of data sparsity in the domain of metaphor detection, the introduction of knowledge distillation serves to prevent over-confidence in the model, thereby further enhancing its ability to detect metaphors.
@@
${{\mathcal{L}}_{s}}=\alpha\mathcal{L}_{GT}^{S}+(1-\alpha){{\tau}^{2}}{{\mathcal{L}}_{KD}}$ (2)
Where $\alpha$ is the balance coefficient. Where $\mathcal{L}_{GT}^{S}$ is the ground-truth loss obtained using one-hot labels, and ${{\mathcal{L}}_{KD}}$ is the knowledge distillation loss obtained using Kullback-Leibler divergenceHershey and Olsen (2007). $\tau$ is the temperature coefficient, which can alleviate the class imbalance problem and help to narrow the gap between the teacher model and the student model. They are defined as follows:
$\mathcal{L}_{GT}^{S}=-\frac{1}{|K|}\sum\limits_{i=1}^{K}{{{y}_{i}}\log\frac{\exp({{e}_{i}})}{\sum\nolimits_{k=1}^{\text{ }\!\!|\!\!\text{ }{{\text{V}}_{a}}\text{ }\!\!|\!\!\text{ }}{\exp({{e}_{k}})}}}$ (3)
${{\mathcal{L}}_{KD}}=\sum\limits_{i=1}^{K}{{{{\tilde{P}}}_{T}}(i)\log\frac{{{{\tilde{P}}}_{T}}(i)}{{{{\tilde{P}}}_{S}}(i)}}$ (4)
@@ 4.1 Datasets
VUA ALLtr 116,622 11.19 6,323 18.4
VUA ALLdev 38,628 11.62 1,550 24.9
VUA ALLte 50,175 12.44 2,694 18.6
VUA Verbtr 15,516 27.90 7,479 20.2
VUA Verbdev 1,724 26.91 1,541 25.0
VUA Verbte 5,783 29.98 2,694 18.6
@@ 4.3 Experimental Settings
In our experiments, we utilize the RoBERTa model Liu et al. (2019) provided by HuggingFace as the encoder. This model consists of 12 layers of Transformer encoders, each comprising 12 attention heads. The hidden dimension is set to 768 for each layer, including both MIP and SPV layers. For knowledge distillation, we employ the MisNet model Zhang and Liu (2022) as the teacher model.
VUA ALL: For VUA ALL dataset, the learning rate is 1e-5 with learning rate warmup, the number of epochs is 10, and the batch size is 64. VUA Verb: For VUA Verb, the learning rate is 1e-5 and the learning rate warmup is used, the number of epochs is 6, and the batch size is 64. MOH-X: For MOH-X, the learning rate is fixed to 1e-5, the number of epochs is 15, and the batch size is 64. TroFi: For TroFi, we only use it for zero-shot evaluation. Where the model is trained on VUA ALL and tested on TroFi. We performed experiments for five times and reported the results with the average. All experiments were done on one NVIDIA RTX 3090 GPU. Our code, saved model weights, and dataset are available on the website.
@@ 5.3 Ablation Study
Figure 4: Mean Acc./F1 scores of different hyparameters on VUA Verb dataset.
5.4 Influence of Hyperparameter in KD
@@ 5.5 Analysis on Knowledge Distillation
Following the introduction of knowledge distillation, the student model adeptly assimilates the knowledge from the teacher model, resulting in rapid enhancement of its capabilities. As illustrated in Figure 5, in the absence of knowledge distillation, the model struggles to acquire meaningful features during the initial stages of training. Moreover, exacerbated by the severe data sparsity issue, the model tends to exhibit over-confidence, leading to disproportionately low F1 scores despite high accuracy rates.

### v2

# Enhancing Metaphor Detection through Soft Labels and Target Word Prediction

@@ Abstract
Metaphors play a significant role in our everyday communication, yet detecting them presents a challenge. Traditional methods often struggle with improper application of language rules and a tendency to overlook data sparsity. To address these issues, we integrate knowledge distillation and prompt learning into metaphor detection. Our approach revolves around a tailored prompt learning framework specifically designed for metaphor detection. By strategically masking target words and providing relevant prompt data, we guide the model to accurately predict the contextual meanings of these words. This approach not only mitigates confusion stemming from the literal meanings of the words but also ensures effective application of language rules for metaphor detection. Furthermore, we’ve introduced a teacher model to generate valuable soft labels. These soft labels provide a similar effect to label smoothing and help prevent the model from becoming over confident and effectively addresses the challenge of data sparsity. Experimental results demonstrate that our model has achieved state-of-the-art performance, as evidenced by its remarkable results across various datasets.
Enhancing Metaphor Detection through Soft Labels and Target Word Prediction
@@ 1 Introduction
The task of metaphor detection poses significant challenges. Early methods Turney et al. (2011); Broadwell et al. (2013); Tsvetkov et al. (2014); Bulat et al. (2017) relied on hand-designed linguistic features to identify metaphors, while later approaches Wu et al. (2018); Gao et al. (2018); Mao et al. (2019) utilized recurrent neural networks (RNNs) to analyze contextual information. However, both categories heavily depended on crafted structures, and the encoders employed struggled to handle complex contextual information, leading to limited effectiveness. In contrast, more recent methods Gong et al. (2020); Su et al. (2020); Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023) leverage pre-trained language models such as BERT Devlin et al. (2019) or RoBERTa Liu et al. (2019) to process contextual information, resulting in significantly improved performance.
While utilizing pre-trained language models has yielded the best results in metaphor detection, these approaches often fail to fully leverage linguistic rules. Among the most advanced methods Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023), the incorporation of linguistic rules, such as the Metaphor Identification Procedure (MIP) Pragglejaz Group (2007), stands out. MIP operates on the principle of extracting the contextual meaning of a target word in the original sentence. If this contextual meaning diverges from the literal meaning, a metaphorical use is identified. For instance, in the sentence "We must bridge the gap between employees and management," the contextual meaning of the target word ’bridge’ pertains to reducing differences or divisions, while its literal meaning refers to constructing a physical bridge. The inconsistency between these meanings indicates a metaphorical use. However, despite the logical validity of MIP, current methods fail to effectively utilize it. Typically, these methods Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023) encode the sentence directly using an encoder and extract the word vector from the position of the target word to represent its contextual meaning. Yet, this approach is flawed as the word vector may be influenced by the literal meaning of the target word, leading to an incomplete representation of its contextual meaning.
@@ 3.2 Metaphor Detection
In this paper, we leverage two linguistic rules to aid in metaphor detection: Selectional Preference Violation (SPV) Wilks (1978) and Metaphor Identification Procedure (MIP) Pragglejaz Group (2007). SPV operates on the principle of identifying metaphors by discerning inconsistencies between the target word and its context. Conversely, MIP detects metaphors by comparing the contextual meaning of the target word with its literal meaning.
When employing SPV linguistic rules for metaphor detection, we adopt a methodology similar to previous works Choi et al. (2021); Zhang and Liu (2022); Zhou et al. (2023). Specifically, We input the sentence containing the target word into an encoder, extract the sentence vector and the target word vector, and input the two vectors as inputs into the full connection layer, through the full connection layer to extract the relationship between the sentence and the target word. Note that the target word here is not replaced with the word generated by the prompt learning.
@@
As for the extraction of the literal meaning of the target word, we adopt the method proposed in the previous work Zhang and Liu (2022) to find out the sentence corresponding to its literal meaning usage for each target word, and put the sentence into the encoder to extract the vector of the position of the target word as the literal meaning of the target word.
@@ 3.3 Knowledge Distillation
Our proposed knowledge distillation method comprises a teacher model T and a student model S. The teacher model T is pre-trained with extensive prior knowledge, enabling it to generate meaningful soft labels to facilitate the optimization process of the student model, and here we use MisNet Zhang and Liu (2022) as the teacher model. By learning from these soft labels generated by the teacher model, the student model S rapidly enhances its capabilities. Furthermore, given the prevalent issue of data sparsity in the domain of metaphor detection, the introduction of knowledge distillation serves to prevent over-confidence in the model, thereby further enhancing its ability to detect metaphors.
@@
${{\mathcal{L}}_{s}}=\alpha\mathcal{L}_{hard}+(1-\alpha){{\mathcal{L}}_{soft}}$ (2)
Where $\alpha$ is the balance coefficient. Where $\mathcal{L}_{hard}$ is the ground-truth loss obtained using one-hot hard labels, and ${{\mathcal{L}}_{soft}}$ is the knowledge distillation loss obtained using soft labels. They are defined as follows:
$\mathcal{L}_{hard}=-\frac{1}{|N|}\sum\limits_{i=1}^{N}{{{y}_{i}}\log({\hat{y_{i}}})}$ (3)
${{\mathcal{L}}_{soft}}=-\frac{1}{|N|}\sum\limits_{i=1}^{N}{{{p}_{i}}\log({{q_{i}}})}$ (4)
@@ 4.1 Datasets
[…]
@@ 4.3 Experimental Settings
In our experiments, we utilize the RoBERTa model Liu et al. (2019) provided by HuggingFace as the encoder. For knowledge distillation, we employ the MisNet model Zhang and Liu (2022) as the teacher model.
VUA ALL: For VUA ALL dataset, the learning rate is 1e-5 with learning rate warmup, the number of epochs is 10, and the batch size is 64. VUA Verb: For VUA Verb, the learning rate is 1e-5 and the learning rate warmup is used, the number of epochs is 6, and the batch size is 64. MOH-X: For MOH-X, the learning rate is fixed to 1e-5, the number of epochs is 15, and the batch size is 64. TroFi: For TroFi, we only use it for zero-shot evaluation. Where the model is trained on VUA ALL and tested on TroFi.
@@ 5.4 Analysis on Knowledge Distillation
Following the introduction of knowledge distillation, the student model adeptly assimilates the knowledge from the teacher model, resulting in rapid enhancement of its capabilities. As illustrated in Fig. 4, in the absence of knowledge distillation, the model struggles to acquire meaningful features during the initial stages of training. Moreover, exacerbated by the severe data sparsity issue, the model tends to exhibit over-confidence, leading to disproportionately low F1 scores despite high accuracy rates.
@@
In addition, we have studied the effects of parameters in knowledge distillation, more details can be found in the appendix A.
5.5 Analysis on Prompt Learning
While the incorporation of prompt learning enhances the precision with which the contextual significance of a target word is represented, there is a potential risk of substituting the target word with one that bears no relevance to the source material, thereby undermining the model’s predictive capabilities. As illustrated in Table 4, the elimination of prompt learning results in a decline in the model’s precision, accompanied by a corresponding increase in recall. This suggests that prompt learning primarily refines the model’s accuracy in forecasting positive classes, without necessarily facilitating the model’s ability to misclassify negative instances as positive. Consequently, prompt learning is not inclined to treat the target word as a context-independent entity.